### v1

@@ 3.1 Design of the radar beamformers and RIS response
$\quad+\sum_{\bm{\theta}\in\Theta_{\rm ev}}g^{\prime}\left(\text{BP}_{\rm{tx}}\left(\bm{\theta};\bm{\omega},\{\tilde{\bm{f}}_{q,r}\}_{q=1}^{N_{\rm{sub}}}\right)-\varepsilon_{\rm ev}\right)$
×$\quad\times\sum_{q=1}^{N_{\rm sub}}\bm{B}_{q}(\bm{\theta};\bm{\omega})\bm{\omega}+\sum_{\bm{\theta}\in\Theta_{\rm ja}}g^{\prime}\left(\text{BP}_{\rm{rx}}\left(\bm{\theta};\bm{\omega}\right)-\varepsilon_{\rm ja}\right)$
×$\quad\times\sum_{q=1}^{N_{\rm sub}}\bm{A}_{q}(\bm{\theta};\bm{\omega})\bm{\omega}\Bigg{]},$ (14)
where $g^{\prime}(\chi)=2\max\{0,\chi\}$ is the derivative of $g(\chi)$, and
@@
$\bm{B}_{q}(\bm{\theta};\bm{\omega})$ $=\operatorname{diag}\{\bm{t}_{q}^{*}(\bm{\theta})\}\bm{G}_{q,\rm rx}^{*}\bm{F}_{q}^{*}\operatorname{diag}\{\bm{\gamma}_{q}\}$
×$\quad\times\bm{F}_{q}^{\mathsf{T}}\bm{G}_{q,\rm rx}^{\mathsf{T}}\operatorname{diag}\{\bm{t}_{q}(\bm{\theta})\}.$ (15b)
@@
$\bm{\omega}^{(i)}=\Pi\left(\bm{\omega}^{(i-1)}+\lambda^{(i)}\bm{\nabla}_{\rm ris}^{(i-1)}\right),$ (16)
where $\lambda^{(i)}$ is the step size at the $i$-th iteration (it can be computed, e.g., though a backtracking Armijo line search [41]), and $\Pi(\,\cdot\,)$ is the projection on $\Omega^{D_{\rm ris}}$, i.e., $\Pi(\bm{\omega})=\hat{\bm{\omega}}$, with $\hat{\omega}_{d}=\omega_{d}/|\omega_{d}|$, for [^4] If $\Omega$ is a finite discrete set, then the projection is $\hat{\omega}_{d}=\operatorname*{\arg\min}_{\chi\in\Omega}|\chi-\omega_{d}|$, for $d=1,\ldots,D_{\rm ris}$. The complex unit circle is a good approximation when the number of the RIS phase states is greater than or equal to 8 [42]. $d=1,\ldots,D_{\rm ris}$.
As to the radar beamformers $\{\tilde{\bm{f}}_{q,r}\}_{q=1}^{N_{\rm{sub}}}$, a suboptimal solution can be found by resorting again to the projected gradient method. Specifically, let $\bm{\nabla}_{r,q}$ be the vector of partial derivatives of Problem (3.1) with respect to $\tilde{\bm{f}}_{q,r}$, then we have
$\bm{\nabla}_{r,q}$ $=2\sum_{\bm{\theta}\in\Theta_{i}}\text{BP}_{q,\text{rx}}(\bm{\theta};\bm{\omega})\bm{b}_{q}(\bm{\theta};\bm{\omega})\bm{b}_{q}^{H}(\bm{\theta};\bm{\omega})\bm{U}_{q}\tilde{\bm{f}}_{q,r}$
$\quad-2\beta\Bigg{[}\sum_{\bm{\theta}\in\Theta_{i,{\rm sl}}}g^{\prime}\left(\text{BP}\left(\bm{\theta};\bm{\omega},\{\tilde{\bm{f}}_{q,r}\}_{q=1}^{N_{\rm{sub}}}\right)-\varepsilon_{\rm sl}\right)$
×$\quad\times\text{BP}_{q,\text{rx}}(\bm{\theta};\bm{\omega})\bm{b}_{q}(\bm{\theta};\bm{\omega})\bm{b}_{q}^{H}(\bm{\theta};\bm{\omega})\bm{U}_{q}\tilde{\bm{f}}_{q,r}$
$\quad+\sum_{\bm{\theta}\in\Theta_{\rm ev}}g^{\prime}\left(\text{BP}_{\rm{tx}}\left(\bm{\theta};\bm{\omega},\{\tilde{\bm{f}}_{q,r}\}_{q=1}^{N_{\rm{sub}}}\right)-\varepsilon_{\rm ev}\right)$
×$\quad\times\bm{b}_{q}(\bm{\theta};\bm{\omega})\bm{b}_{q}^{H}(\bm{\theta};\bm{\omega})\bm{U}_{q}\tilde{\bm{f}}_{q,r}\Bigg{]},$ (17)
where
$\bm{b}_{q}(\bm{\theta};\bm{\omega})=\sqrt{\gamma_{q,r}}\bm{G}_{q,\rm tx}^{H}\operatorname{diag}\{\bm{t}_{q}^{*}(\bm{\theta})\}\bm{\omega}^{*}\in\mathbb{C}^{D_{\rm{tx}}}.$ (18)
@@
$\tilde{\bm{f}}_{q,r}^{(i)}=\frac{\tilde{\bm{f}}_{q,r}^{(i-1)}+\lambda^{(i)}\bm{\nabla}_{r,q}^{(i-1)}}{\bigl{\|}\bm{U}_{q}\bigl{(}\tilde{\bm{f}}_{q,r}^{(i-1)}+\lambda^{(i)}\bm{\nabla}_{r,q}^{(i-1)}\bigr{)}\bigr{\|}},$ (19)
@@
Alg. 1 summarizes the proposed block-coordinate ascent routine. Since the projected gradient method starts from the point found at the previous iteration, the objective function is non-decreasing in successive iterations, and Alg. 1 converges. However, since Problem (3.1) is not convex and since the feasible search set cannot be expressed as the Cartesian product of closed convex sets, there is no guarantee that a global maximum is reached.
Algorithm 1 Suboptimal solution to Problem (3.1)
@@
6:        $\bm{\omega}=\bm{\omega}+\lambda\bm{\nabla}_{\rm ris}$
@@
10:        compute $\bm{\nabla}_{r,q}$ in (3.1), $\forall q$
@@
12:        $\tilde{\bm{f}}_{q,r}=\tilde{\bm{f}}_{q,r}+\lambda\bm{\nabla}_{r,q}$, $\forall q$
@@ 3.2 Design of the radar detector
The measurements collected in each CPI are processed through a bank of correlators, each matched to a delay, say $\tau$, and a Doppler shift, say $\nu$, taken from the uniformly-spaced grids $\mathcal{G}_{\rm del}=\{\bar{\tau}_{1},\ldots,\bar{\tau}_{N_{\rm del}}\}$ and $\mathcal{G}_{\rm dop}=\{\bar{\nu}_{1},\ldots,\bar{\nu}_{N_{\rm dop}}\}$, respectively; each correlator is followed by a square law detector and a noise variance normalization. For any inspected triplet $(\bm{\theta},\tau,\nu)\in\mathcal{G}_{\rm dir}\times\mathcal{G}_{\rm del}\times\mathcal{G}_{\rm dop}$, the output statistic is [39]
$A(\bm{\theta},\tau,\nu)=\frac{1}{\sigma_{r}^{2}}\left|\sum_{q=1}^{N_{\rm sub}}\sum_{n=1}^{N_{\rm sym}}\bm{u}_{q}^{\mathsf{H}}(n;\bm{\theta},\tau,\nu)\bm{y}_{q,r}(n)\right|^{2},$ (20)
@@
and the set of $\{w_{q}(n)\}$ are positive weights normalized to have $\sum_{q=1}^{N_{\rm sub}}\sum_{n=1}^{N_{\rm sym}}w^{2}_{q}(n)=1$. In (21), the knowledge of the transmitted signal is exploited to combine the measurements (and coherently integrate the energy) taken in different subcarriers and symbol intervals; also, the adopted normalization ensures that $\|\bm{u}_{q}(n;\bm{\theta},\tau,\nu)\|=1$; finally, the weights $\{w_{q}(n)\}$ can be chosen to obtain lower sidelobes in the range and Doppler domains at the price of a wider and lower mainlobe.
@@
In a standard radar detector, the decision on the target presence is made on the basis of the current scan only. However, this approach shows severe limitations when the signal amplitude is weak compared to the background noise, which can be the case for non-line-of-sight mono-static radar architectures. To overcome this limitation, we propose here to make a decision on the target presence after integrating its signal returns over multiple consecutive scans. This is challenging in the presence of target motion, and TBD techniques [^5] This terminology is used to mark the difference with respect to the classical approach, where the target is first detected and then its motion during multiple scans is tracked; with the TBD approach, instead, a number of scans are jointly processed with no (or little) thresholding taking place, and an estimated trajectory is returned at the same time as a detection is declared. are required to properly integrate the echoes over consecutive scans [19]. Many approaches have been proposed in the literature, and here we resort to the multi-frame detector proposed in [22, 23, 24], which adopts a computationally efficient dynamic programming algorithm to compute the required statistics. The considered TBD scheme is reported in Fig 3. The first stage contains the detector and plot-extractor, wherein the threshold is lowered to obtain a richer set of plots: this allows to preserve the measurements from dim targets at the price of increasing the number of false alarms. The second stage is the TBD processor, which exploits the space-time correlation among the plots in the current scan and those in the previous $N_{\rm scan}-1$ scans to confirm or delete them; the final decision on the target presence in the current scan is then made by comparing a convenient scoring metric against a threshold that is adjusted to have the desired probability of false alarm ($P_{\rm fa}$).
Figure 3: Block diagram of the considered radar detector.
@@
It is worth mentioning that this approach can readily be generalized to the case where multiple targets can be present in the inspected volume $\mathcal{V}_{r}$: data association problems arising when more plots at the current scan correlate with the same plot at some previous scan can be solved through a successive track cancellation procedure [23]. Also, the trajectory of detected targets can be smoothed by using polynomial regression: this improves the accuracy of the range and azimuth measurements based on the estimated history and, as a by-product, may provide information about the velocity of the targets [24].
@@ 4.1 System setup
The forward link between the $j$-th transmitting element and the $i$-th RIS element on the $q$-th subcarrier is modeled as
@@
Figure 5: Jammer and eavesdropper regions, along with the sets $\Theta_{i}$, $\Theta_{i,{\rm sl}}$, $\Theta_{\rm ev}$, and $\Theta_{\rm ja}$ in Problem (3.1).
Figure 6: Realized (normalized) two-way beampattern for the pointing directions $\{\bar{\bm{\theta}}_{i}\}_{i=1}^{6}$; the 3-dB beamwidth is delimited by the dashed line.
@@ 4.2 Numerical results
Figure 7: Probability of target detection (top plot) and position RMSE (bottom plot) versus communication sum rate for various numbers of users and processed scans.
@@
Figure 8: Communication sum rate (top plot), probability of target detection (middle plot), and RMSE in the position estimation (bottom plot) versus the fraction of power devoted to the radar function for different numbers of users when $N_{\rm{scans}}=15$ scans are processed, and the orthogonality constraint in (11) enforced on the radar beamformers is enabled/disabled.
@@ 5 Conclusions
Future works should account for the presence of estimation errors in the user channels. Also, they could further investigate the potential benefits of resorting to a multi-frame processing when the sensing and communication transmit beamformers are jointly designed, when the BS is mounted on a moving platform, or when multiple BSs cover the same region. In particular, the multi-frame radar processing might also improve the localization accuracy of legacy mobile users, thus helping the construction of more efficient schemes for resource scheduling and for channel estimation and prediction.

### v2

@@ 3.1 Design of the radar beamformers and RIS response
$\quad-2\beta\Bigg{[}\sum_{\bm{\theta}\in\Theta_{\rm ev}}\!\!g^{\prime}\Bigl{(}\text{BP}_{\rm{tx}}(\bm{\theta};\bm{\omega},\{\tilde{\bm{f}}_{q,r}\}_{q=1}^{N_{\rm{sub}}})-\varepsilon_{\rm ev}\Bigr{)}\!\!\sum_{q=1}^{N_{\rm sub}}\!\!\bm{B}_{q}(\bm{\theta};\bm{\omega})\bm{\omega}\Bigg{]}\!\!-\!2\beta\Bigg{[}\sum_{\bm{\theta}\in\Theta_{\rm ja}}\!\!g^{\prime}\Bigl{(}\text{BP}_{\rm{rx}}(\bm{\theta};\bm{\omega})-\varepsilon_{\rm ja}\Bigr{)}\!\!\sum_{q=1}^{N_{\rm sub}}\!\!\bm{A}_{q}(\bm{\theta};\bm{\omega})\bm{\omega}\Bigg{]}\!,$ (14)
@@
$\bm{B}_{q}(\bm{\theta};\bm{\omega})$ $=\operatorname{diag}\{\bm{t}_{q}^{*}(\bm{\theta})\}\bm{G}_{q,\rm tx}^{*}\bm{F}_{q}^{*}\operatorname{diag}\{\bm{\gamma}_{q}\}$
×$\quad\times\bm{F}_{q}^{\mathsf{T}}\bm{G}_{q,\rm tx}^{\mathsf{T}}\operatorname{diag}\{\bm{t}_{q}(\bm{\theta})\}.$ (15b)
@@
$\bm{\omega}^{(i)}=\Pi\left(\bm{\omega}^{(i-1)}+\lambda^{(i)}\bm{\nabla}_{\!\!\bm{\omega}}^{(i-1)}\right),$ (16)
where $\lambda^{(i)}$ is the step size at the $i$-th iteration, which can be computed, for example, though a backtracking Armijo line search [48], and $\Pi(\,\cdot\,)$ is the projection on $\Omega^{D_{\rm ris}}$, i.e., $\Pi(\bm{\omega})=\hat{\bm{\omega}}$, with $\hat{\omega}_{d}=\omega_{d}/|\omega_{d}|$, for [^6] If $\Omega$ is a finite discrete set, then the corresponding projection is given by $\hat{\omega}_{d}=\operatorname*{\arg\min}_{\chi\in\Omega}|\chi-\omega_{d}|$, for $d=1,\ldots,D_{\rm ris}$. The complex unit circle is a good approximation when the number of the RIS phase states is greater than or equal to 8 [49]. $d=1,\ldots,D_{\rm ris}$.
As to the radar beamformers $\{\tilde{\bm{f}}_{q,r}\}_{q=1}^{N_{\rm{sub}}}$, a suboptimal solution can be found by resorting again to the projected gradient method. Specifically, let $\bm{\nabla}_{q,r}$ be the vector of partial derivative of the objective function in Problem (13) with respect to $\tilde{\bm{f}}_{q,r}$, that is reported in (17) at the top of the next page, where
$\bm{\nabla}_{q,r}$ $=2\sum_{\bm{\theta}\in\Theta_{i}}\text{BP}_{q,\text{rx}}(\bm{\theta};\bm{\omega})\bm{b}_{q}(\bm{\theta};\bm{\omega})\bm{b}_{q}^{\mathsf{H}}(\bm{\theta};\bm{\omega})\tilde{\bm{f}}_{q,r}$
$\quad-2\beta\Bigg{[}\sum_{\bm{\theta}\in\Theta_{i,{\rm sl}}}g^{\prime}\Bigl{(}\text{BP}(\bm{\theta};\bm{\omega},\{\tilde{\bm{f}}_{q,r}\}_{q=1}^{N_{\rm{sub}}})-\varepsilon_{\rm sl}\Bigr{)}\text{BP}_{q,\text{rx}}(\bm{\theta};\bm{\omega})\bm{b}_{q}(\bm{\theta};\bm{\omega})\bm{b}_{q}^{\mathsf{H}}(\bm{\theta};\bm{\omega})\tilde{\bm{f}}_{q,r}\Bigg{]}$
$\quad-2\beta\Bigg{[}\sum_{\bm{\theta}\in\Theta_{\rm ev}}g^{\prime}\Bigl{(}\text{BP}_{\rm{tx}}(\bm{\theta};\bm{\omega},\{\tilde{\bm{f}}_{q,r}\}_{q=1}^{N_{\rm{sub}}})-\varepsilon_{\rm ev}\Bigr{)}\bm{b}_{q}(\bm{\theta};\bm{\omega})\bm{b}_{q}^{\mathsf{H}}(\bm{\theta};\bm{\omega})\tilde{\bm{f}}_{q,r}\Bigg{]},$ (17)
$\bm{b}_{q}(\bm{\theta};\bm{\omega})\!=\!\sqrt{\gamma_{q,r}}\bm{U}_{q}^{\mathsf{H}}\bm{G}_{q,\rm tx}^{\mathsf{H}}\operatorname{diag}\{\bm{t}_{q}^{*}(\bm{\theta})\}\bm{\omega}^{*}\in\mathbb{C}^{D_{\rm{tx}}-K}\!\!\!.$ (18)
@@
$\tilde{\bm{f}}_{q,r}^{(i)}=\frac{\tilde{\bm{f}}_{q,r}^{(i-1)}+\lambda^{(i)}\bm{\nabla}_{q,r}^{(i-1)}}{\bigl{\|}\bm{U}_{q}\bigl{(}\tilde{\bm{f}}_{q,r}^{(i-1)}+\lambda^{(i)}\bm{\nabla}_{q,r}^{(i-1)}\bigr{)}\bigr{\|}},$ (19)
@@
Alg. 1 summarizes the proposed block-coordinate ascent routine. Since the projected gradient method starts from the point found at the previous iteration, the objective function is non-decreasing in successive iterations, and Alg. 1 converges. However, since Problem (13) is not convex and since the feasible search set cannot be expressed as the Cartesian product of closed convex sets, there is no guarantee that a global maximum is reached.
Algorithm 1 Suboptimal solution to Problem (13)
@@
6:        $\bm{\omega}=\bm{\omega}+\lambda\bm{\nabla}_{\!\!\bm{\omega}}$
@@
10:        compute $\bm{\nabla}_{q,r}$ in (17), $\forall q$
@@
12:        $\tilde{\bm{f}}_{q,r}=\tilde{\bm{f}}_{q,r}+\lambda\bm{\nabla}_{q,r}$, $\forall q$
@@ 3.2 Design of the radar detector
The measurements collected in each CPI are processed through a bank of correlators, each matched to a delay, say $\tau$, and a Doppler shift, say $\nu$, taken from the uniformly-spaced grids $\mathcal{G}_{\rm del}=\{\bar{\tau}_{1},\ldots,\bar{\tau}_{N_{\rm del}}\}$ and $\mathcal{G}_{\rm dop}=\{\bar{\nu}_{1},\ldots,\bar{\nu}_{N_{\rm dop}}\}$, respectively; each correlator is followed by a square law detector and a noise variance normalization. Following [46], the output statistic is
$A(\bm{\theta},\tau,\nu)=\frac{1}{\sigma_{r}^{2}}\left|\sum_{q=1}^{N_{\rm sub}}\sum_{n=1}^{N_{\rm sym}}\bm{u}_{q}^{\mathsf{H}}(n;\bm{\theta},\tau,\nu)\bm{y}_{q,r}(n)\right|^{2}$ (20)
@@
for $q=1,\ldots,N_{\rm sub}$ and $n=1,\ldots,N_{\rm sym}$, are the tap amplitudes of the correlator matched to the triplet $(\bm{\theta},\tau,\nu)\in\mathcal{G}_{\rm dir}\times\mathcal{G}_{\rm del}\times\mathcal{G}_{\rm dop}$ and $\{w_{q}(n)\}$ are positive weights normalized to have $\sum_{q=1}^{N_{\rm sub}}\sum_{n=1}^{N_{\rm sym}}w^{2}_{q}(n)=1$. In (21), the knowledge of the transmitted signal is exploited to combine the measurements (and coherently integrate the energy) taken in different subcarriers and symbol intervals; also, the adopted normalization ensures that $\|\bm{u}_{q}(n;\bm{\theta},\tau,\nu)\|=1$; finally, the weights $\{w_{q}(n)\}$ can be chosen to obtain lower sidelobes in the range and Doppler domains at the price of a wider and lower mainlobe.
@@
In a standard radar detector, the decision on the target presence is made on the basis of the current scan only. However, this approach shows severe limitations when the signal amplitude is weak compared to the background noise, which can be the case for non-line-of-sight mono-static radar architectures. To overcome this limitation, we propose here to make a decision on the target presence after integrating its signal returns over multiple consecutive scans. This is challenging in the presence of target motion, and TBD techniques [^7] This terminology is used to mark the difference with respect to the classical approach, where the target is first detected and then tracked over multiple scans; with the TBD approach, instead, a number of scans are jointly processed with no (or little) thresholding taking place, and an estimated trajectory is returned at the same time as a detection is declared. are required to properly integrate the echoes over consecutive scans [24].Many approaches have been proposed in the literature. Here we resort to the multi-frame detector proposed in [27, 28], that has also been validated with real data in [29]. The detector adopts a computationally efficient dynamic programming algorithm to compute the required statistics. The considered TBD scheme is reported in Fig 3. The first stage contains the detector and plot-extractor, wherein the threshold is lowered to obtain a richer set of plots: this allows to preserve the measurements from dim targets at the price of increasing the number of false alarms. The second stage is the TBD processor, which exploits the space-time correlation among the plots in the current scan and those in the previous $N_{\rm scan}-1$ scans to confirm or delete them; the final decision on the target presence in the current scan is then made by comparing a convenient scoring metric against a threshold that is adjusted to have the desired probability of false alarm ($P_{\rm fa}$).
@@
The complexity of this TBD processor is linear in the number of integrated scans, $N_{\rm{scan}}$, and quadratic in the number of plots per scan, i.e, the number of rows in $\bm{S}_{\ell}$ for $\ell=1,\ldots,N_{\rm{scan}}$, which makes its real time implementation feasible: a detailed discussion on this can be found in [27, 28]. Also, it is worth mentioning that the trajectory of the detected target can be smoothed by using polynomial regression: this improves the accuracy of the range and azimuth measurements based on the estimated history and, as a by-product, may provide information about the velocity of the targets [29]. Finally, notice that this method can readily be generalized to the case where multiple targets can be present in the inspected volume $\mathcal{V}_{r}$: data association problems arising when more plots at the current scan correlate with the same plot at some previous scan can be solved through a successive track cancellation procedure, with a complexity at most linear in the number of detected targets [27, 28].
Figure 3: Block diagram of the considered radar detector.
@@ 4.1 System setup
The link between the $j$-th transmitting element and the $i$-th RIS element on the $q$-th subcarrier is modeled as
@@
Figure 5: Jammer and eavesdropper regions, along with the sets $\Theta_{i}$, $\Theta_{i,{\rm sl}}$, $\Theta_{\rm ev}$, and $\Theta_{\rm ja}$ in Problem (12).
Figure 6: Normalized two-way beampattern for the considered pointing directions $\{\bar{\bm{\theta}}_{i}\}_{i=1}^{6}$. In each plot, the inspected, eavesdropper, and jammer regions are also shown, along with a dashed black contour line marking the 3-dB level of the mainlobe.
@@ 4.2 Numerical results
Figure 7: Probability of target detection (top) and RMSE on the position estimation (bottom) versus the user sum-rate for various numbers of users and scans. Here, the number of scans is $N_{\rm{scans}}=1,5,8,12,15$, which increases in the direction of the arrow and corresponds to markers
$\blacklozenge$
$\blacksquare$
[…]
$\bullet$
, respectively.
@@
Figure 8: User sum-rate (top), probability of target detection (middle), and RMSE in the position estimation (bottom) versus the fraction of power devoted to the radar function for different numbers of users when $N_{\rm{scans}}=15$ scans are processed, and the orthogonality constraint in (11) enforced on the radar beamformers is enabled/disabled.
@@ 5 Conclusions
Future works should account for the presence of estimation errors in the user channels. Also, they could further investigate the potential benefits of resorting to a multi-frame processing when the sensing and communication transmit beamformers are jointly designed, when the BS is mounted on a moving platform, or when multiple BSs cover the same region. In particular, the multi-frame radar processing might also improve the localization accuracy of legacy mobile users, thus helping the construction of more efficient schemes for resource scheduling and for channel estimation and prediction. Finally, an experimental validation assessing the achievable performance under real-world operating conditions is still missing and should be verified and discussed in future studies.